# Between Circuits and Chomsky: Pre-pretraining on Formal Languages Imparts Linguistic Biases

### v1

@@ 3.2 Between Circuits and Chomsky
We hypothesize that a good pre-pretraining language should both encapsulate the complexity of natural language and be robustly learnable by transformers in a way that length generalizes. Because natural language is hierarchically structured and C-RASP has been suggested as a formal model of what transformers can learn robustly, this motivates the following hypothesis:
@@
To test this hypothesis, we pre-pretrain language models on the following formal languages:
@@
    1-Dyck: the nested parentheses language. Context-free, in C-RASP.
@@
[…]
@@
    $k$-Shuffle Dyck: $k$-Dyck with cross-serial dependencies. Context-sensitive, in C-RASP.
@@
[…]
The three variants of Dyck languages model hierarchical structure, while $ww$, like the act of copying, has a fixed dependency structure. For examples of these languages, see Table 1. Full proofs of where these languages lie on the Chomsky and circuit hierarchies can be found in Appendix A. [^3] The languages NEST and CROSS from Papadimitriou and Jurafsky (2023) are instances of $k$-Dyck and $k$-Shuffle Dyck, respectively. Their results align with our hypothesis.
@@
Figure 2: Evaluating models at the optimal amount of pre-pretraining $t_{0}^{*}$ for each formal language. $k$-Shuffle Dyck performs the best overall on our evaluation metrics.
We deliberately chose languages that are similar to each other. $k$-Dyck and $k$-Shuffle Dyck can be seen as different extensions of 1-Dyck: $k$-Dyck swaps out paired parentheses in valid 1-Dyck strings with new parentheses pairs, while $k$-Shuffle Dyck effectively interleaves several 1-Dyck sequences Suzgun et al. (2019). Finally, $ww$ contrasts with $k$-Shuffle Dyck as a maximally context-sensitive language, since all the dependencies in $ww$ are cross-serial.
To reduce confounds, we construct 1-Dyck, $k$-Dyck, and $k$-Shuffle Dyck corpora with matching depth distributions. We do so during corpus generation by randomly opening or closing parentheses with probability $p=0.5$, which yields a harmonic distribution over depths. Similarly, the $k$-Dyck, $k$-Shuffle Dyck, and $ww$ corpora each have 128 unique tokens, or 64 unique parentheses pairs ($k=64$) for Dyck. All language models are pre-pretrained on the same number of tokens with sequence packing.
@@ 4 Testing the Expressivity Hypothesis
For natural language ($\mathcal{D}_{\text{pt}}$), we trained Pythia 160M models Biderman et al. (2023) for 10,000 steps, or roughly 665 million tokens. We use C4 as the natural language dataset Raffel et al. (2019). All training hyperparameter details can be found in Appendix B. For BLiMP accuracies by category, see Figure 9.
Figure 3: Validation loss on C4 as a function of pre-pretraining tokens. For the formal languages that improve validation loss over no pre-pretraining, performance is u-shaped with respect to pre-pretraining tokens, indicating that pre-pretraining is harmful after a certain point.
@@
The most efficient formal language is $k$-Shuffle Dyck, with $t_{0}^{*}=500$. Both $k$-Shuffle Dyck and $k$-Dyck outperform natural language pre-pretraining, but $k$-Dyck’s optimum is achieved around $t_{0}^{*}=1000$. $ww$ pre-pretraining is unhelpful at all durations. The other languages, which positively transfer to natural language, exhibit a roughly u-shaped scaling pattern: after reaching the optimal training duration, training more on formal language is harmful for natural language learning. Shuffle Dyck has the highest MRS, indicating that it replaces tokens on natural language most efficiently. Furthermore, the MRS for 1-Dyck, $k$-Dyck, and $k$-Shuffle Dyck are all greater than 1, indicating that exchanging natural language for these formal languages is compute-optimal in our setting. See Table 5 in the Appendix for MRS values.
@@
$k$-Shuffle Dyck is the best-performing formal language on the validation set of C4, followed by $k$-Dyck (Figure 2). Interestingly, pre-pretraining on all four formal languages improves accuracy in grammaticality, but pre-pretraining on natural language (C4) does not. This indicates that formal language pre-pretraining also changes models’ generalization properties, in addition to driving the language modeling loss lower. We hypothesize this is because pre-pretraining induces representations useful for modeling hierarchical structure, and we revisit this point in §5.
@@
5 Pruning
@@
Subnetworks hypothesis: Subnetworks established during formal language pre-pretraining are later used to represent the hierarchical structure of natural language.
We test this hypothesis by ablating attention heads of the pre-pretraining subnetwork and comparing the drop in performance against random attention head ablations. Concretely, we pre-pretrain on $\mathcal{D}_{\text{ppt}}$ and prune the model to find the sparse subnetwork $\mathcal{M}(\theta_{t_{0}})$. We use the heuristic core pruning algorithm from Bhaskar et al. (2024), which iteratively removes attention heads from the transformer using structured pruning (Xia et al., 2022) while minimizing the tradeoff between sparsity and language modeling loss on $\mathcal{D}_{\text{ppt}}$. After transfer and training on $\mathcal{D}_{\text{pt}}$, we evaluate the masked model $\mathcal{M}(\theta_{t_{1}})$ against networks with the same number of attention heads, randomly ablated $\left(\mathcal{M}_{\text{null}}(\theta_{t_{1}})\right)$.
Positive transfer from $\mathcal{D}_{\text{ppt}}$ to $\mathcal{D}_{\text{pt}}$ could occur for reasons unrelated to subnetworks. In this case, the masked model $\mathcal{M}$ should perform no better than random masks $\mathcal{M}_{\text{null}}$ when applied to $\theta_{t_{1}}$. However, if pre-pretraining does induce useful inductive biases, we would expect $\mathcal{M}$ to be an important subnetwork even after training on $\mathcal{D}_{\text{pt}}$. So in the alternative hypothesis, $\mathcal{M}$ should significantly outperform $\mathcal{M}_{\text{null}}$ in validation loss and zero-shot tasks.
@@
Results.
After pre-pretraining on $k$-Shuffle Dyck, we ablate $50\%$ of the attention heads using structured pruning to create the sparse subnetwork $\mathcal{M}$. Following previous work Bhaskar et al. (2024); Zhang and Nanda (2024), we replace an ablated head with its mean activation. In addition to random masked subnetworks $\mathcal{M}_{\text{null}}$, we also compare $\mathcal{M}$ against its complement subnetwork $\mathcal{M}^{c}$ on our evaluation metrics.
@@ Results.
In Figure 4, we find that $\mathcal{M}$ outperforms random masking and $\mathcal{M}^{c}$. We reject the null hypothesis that the subnetwork $\mathcal{M}$ established during pre-pretraining has the same performance as a randomly sampled subnetwork ($p\ll 0.001)$. With $\mathcal{M}^{c}$ performing poorly, this suggests that the subnetwork learned during pre-pretraining is indeed being reused in natural language modeling. We also note that $\mathcal{M}$ does not achieve the performance of the full network, indicating that attention heads outside of $\mathcal{M}$ are also useful for natural language.
@@
Figure 5: Pre-pretraining on $n$-gram metamers Kumar et al. (2022) of $k$-Shuffle Dyck performs worse than pre-pretraining on $k$-Shuffle Dyck itself. The best vocabulary size for $k$-Shuffle Dyck we observe is $k=128$.
Figure 6: Pre-pretraining Pythia-1B on $k$-Shuffle Dyck continues to outperform baselines, even after 1.6B pretraining tokens.
A Syntactic Subnetwork.
Last, we provide evidence that the subnetwork $\mathcal{M}$ mostly implements syntax, but that pre-pretraining also helps models learn non-syntactic knowledge like morphology, or word structure. To observe this, we first compare $\mathcal{M}$’s performance on BLiMP against that of the full network, aggregated by the classification of linguistic phenomena as reflecting morphology (27% of BLiMP examples) versus syntax and semantics (73%). In Figure 7 (left), we see that morphology declines in performance from full to $\mathcal{M}$ much more than syntax and semantics. This suggests the attention heads driving performance on morphology lie outside of $\mathcal{M}$, aligning with the intuition that pre-pretraining on formal languages should form a subnetwork attending mainly to structure. However, when we examine the performance gain over no pre-pretraining (Figure 7 right), we find that morphology also benefits from pre-pretraining. We conclude that pre-pretraining has second-order effects. In addition to syntactic benefits, pre-pretraining can improve performance on non-syntactic tasks.
However, pre-pretraining is also not a free lunch. Figure 9, which shows BLiMP accuracies by subtask, includes some subtasks where no pre-pretraining outperforms $k$-Shuffle Dyck pre-pretraining. Thus, pre-pretraining, mediated through the subnetwork $\mathcal{M}$, can also hinder performance. Understanding how to tailor the improvements of pre-pretraining to one’s use case is an exciting area for future work.
Figure 7: Pruning the model to $\mathcal{M}$ decreases morphological performance more than performance on syntax and semantics in BLiMP (left), but pre-pretraining itself still has a positive impact on morphological performance (right).
6 Analyzing Shuffle Dyck
In §4, we observed that pre-pretraining on $k$-Shuffle Dyck helps models learn natural language. We now perform some additional analyses that help qualify this result. First, we examine whether the rule-based structure of $k$-Shuffle Dyck is truly necessary, over a facsimile dataset with the same statistical properties. Next, we study the impact that varying the vocabulary size $k$ of $k$-Shuffle Dyck has on performance. Finally, we perform a larger scale training run with Pythia 1B and find that pre-pretraining on $k$-Shuffle Dyck helps in this setting as well. Results for this section are in Figures 6 and 6.
Order statistics.
@@ 6 Analyzing Shuffle Dyck
We seek to understand whether some low-level statistical properties of $k$-Shuffle Dyck explain its success, instead of its dependency structure. Indeed, neural networks can exhibit distributional simplicity bias (DSB)—they learn simpler statistical patterns, such as the mean and covariance of their representations, before progressing to higher-order relationships Saxe et al. (2013); LeCun et al. (1991). Belrose et al. (2024) also demonstrate this phenomenon in transformer language models, which learn $n$-gram statistics in order of increasing complexity.
@@
Accordingly, we ablate the rule-based structure of $k$-Shuffle Dyck by training unigram, bigram, and trigram models on our pre-pretraining corpus. We then generate from these $n$-gram models to produce metamer datasets Kumar et al. (2022) of equivalent size. These metamers match $k$-Shuffle Dyck up to some $n$-gram statistic, but do not capture its rule.
We find that pre-pretraining on metamer datasets is strictly worse than pre-pretraining on $k$-Shuffle Dyck. However, the results follow a predictable pattern—pre-pretraining on the unigram metamer performs the worst, followed by bigram and trigram. This indicates that DSB might also apply within pre-pretraining.
Vocabulary size.
@@
Larger scale.
Finally, we examine whether our results generalize to larger settings by training Pythia-1B for 1.63B tokens on C4 (25,000 steps). In this setting, pre-pretraining on $k$-Shuffle Dyck continues to outperform on all evaluation metrics (Figure 6) and achieves the final loss of no pre-pretraining in 1.10B total tokens. This equates to a token efficiency gain of 33% $\left(1-\frac{1.10\text{B}}{1.63\text{B}}\right)$, or an MRS of 17.3 $\left(\frac{1.63\text{B}-1.10\text{B}}{0.03\text{B}}\right)$. $k$-Shuffle Dyck’s MRS is $\gg$1 for both 160M and 1B training runs, suggesting that pre-pretraining could increase the efficiency of large-scale pretraining as well.
@@ 7 Related Work
Efficient Learning.
The goal of pre-pretraining is similar to that of optimization-based meta-learning, which aims to create a weight initialization that allows the model to rapidly learn new tasks Finn et al. (2017); Nichol et al. (2018). McCoy and Griffiths (2023) use meta-learning over procedurally-generated context-free languages to rapidly learn formal linguistic patterns. Positive transfer in NLP also occurs across both domains and languages (Ruder et al., 2019; Pruksachatkun et al., 2020; Deshpande et al., 2022).
Linguistic Generalization.
In addition to formal languages, various other datasets can also change the generalization properties of neural networks. Mueller and Linzen (2023) show that pretraining on child-directed speech gives a better inductive bias for learning hierarchical syntactic features than standard pretraining corpora. Furthermore, introducing a small amount of synthetic, disambiguating data into pretraining can induce a language model to change its generalization strategy Warstadt et al. (2020b). Algorithmic synthetic data (sampled from a formal language) is often used to test the generalization properties of neural networks McCoy et al. (2019); Kim and Linzen (2020); Li et al. (2023).
Curriculum Learning.
Pre-pretraining can be seen as curriculum learning (Bengio et al., 2009), where data increases in difficulty during training. Many language model training recipes have several different stages of data mixtures Allal et al. (2025); OLMo et al. (2025); Ouyang et al. (2022). Recent work has also attempted to create algorithms that automate the discovery of language modeling curricula Chen et al. (2025); Jiang et al. (2025).

### v2

@@ 3.2 Between Circuits and Chomsky
We hypothesize that a good pre-pretraining language should both mimic particular aspects of the complexity of natural language and be robustly learnable by transformers in a way that generalizes to longer strings than observed in training. Because, as we discussed in §2, natural language is hierarchically structured and C-RASP is a formal model of what transformers can learn robustly, this motivates the following hypothesis:
@@
To test this hypothesis, we pre-pretrain transformer language models on the following formal languages:
@@
    1-Dyck: the nested parentheses language. This language is context-free and in C-RASP.
@@
    $k$-Dyck: contains $k$ different types of parentheses. The language is context-free and in FO(M) but not in C-RASP.
@@
    $k$-Shuffle Dyck: $k$-Dyck with cross-serial dependencies. This language is context-sensitive and in C-RASP. [^4] Figure 10 shows a minimal code snippet for C-RASP.
@@
    $ww$: The copy language. This language is context-sensitive and in FO(M) but not in C-RASP.
@@
The three variants of Dyck languages model hierarchical structure, while $ww$ has a fixed dependency structure that maps the first half of the string onto the second half (Table 1). Proofs of where these languages lie on the Chomsky and circuit hierarchies can be found in Appendix A. [^5] The languages NEST and CROSS from Papadimitriou and Jurafsky (2023) are instances of $k$-Dyck and $k$-Shuffle Dyck, respectively. Their results align with our hypothesis.
Figure 2: Evaluating models on overall language modeling loss, grammaticality and retrieval, at the optimal amount of pre-pretraining $t_{0}^{*}$ for each formal language (the value of $t_{0}^{*}$ for each language is determined based on Figure 3).
We deliberately chose languages that are similar to each other. $k$-Dyck and $k$-Shuffle Dyck can be seen as different extensions of 1-Dyck: $k$-Dyck swaps out paired parentheses in valid 1-Dyck strings with new parentheses pairs, while $k$-Shuffle Dyck effectively interleaves several 1-Dyck sequences Suzgun et al. (2019). Finally, $ww$ contrasts with $k$-Shuffle Dyck as a maximally context-sensitive language, since all the dependencies in $ww$ are cross-serial (i.e. none are nested within one another).
We construct 1-Dyck, $k$-Dyck, and $k$-Shuffle Dyck corpora with matching depth distributions by randomly opening or closing parentheses with probability $p=0.5$, which yields a harmonic distribution over depths. We truncate the length of the sequences at 2048. We also match the vocabulary size: $k$-Dyck, $k$-Shuffle Dyck, and $ww$ corpora each have 128 unique vocabulary items, or 64 unique parentheses pairs ($k=64$) for the Dyck languages (we explore the effect of this hyperparameter in §6). All models are pre-pretrained on the same number of tokens with sequence packing.
@@ 4 Testing the Expressivity Hypothesis
Figure 3: C4 validation loss as a function of pre-pretraining tokens. For the formal languages that improve validation loss over no pre-pretraining, there is an optimal training duration after which additional pre-pretraining is harmful.
For natural language ($\mathcal{D}_{\text{pt}}$), we trained Pythia 160M models Biderman et al. (2023) for 10,000 steps, or roughly 665 million tokens. We use C4 as the natural language dataset Raffel et al. (2019). For training hyperparameters, see Appendix B.
@@
While both $k$-Shuffle Dyck and $k$-Dyck outperform natural language pre-pretraining, $k$-Shuffle Dyck is more efficient with $t_{0}^{*}=500$ compared to $t_{0}^{*}=1000$ for $k$-Dyck. Pre-pretraining on $ww$ is unhelpful at all durations. For each of the languages where pre-pretraining is effective, there is an optimal duration after which additional formal language pre-pretraining leads to less effective transfer overall. $k$-Shuffle Dyck has the highest MRS, indicating that it replaces tokens on natural language most efficiently (see Table 5 in the Appendix). Furthermore, the MRS for 1-Dyck, $k$-Dyck, and $k$-Shuffle Dyck are all greater than 1, indicating that exchanging natural language for these formal languages is compute-optimal in our setting.
@@
$k$-Shuffle Dyck is the best-performing formal language on the natural language validation set from C4, followed by $k$-Dyck (Figure 2). Interestingly, pre-pretraining on all four formal languages improves accuracy in grammaticality, but pre-pretraining on natural language does not (for grammaticality accuracies by category, see Figure 8 in the Appendix). This indicates that formal language pre-pretraining also changes models’ generalization properties, in addition to driving the language modeling loss lower. We hypothesize this is because pre-pretraining induces representations useful for modeling hierarchical structure; we revisit this point in §5.
@@
Summary.
Hierarchical dependencies, which both $k$-Dyck and $k$-Shuffle Dyck have, appear to be crucial for positive transfer from formal to natural language. Although of these two languages only $k$-Shuffle Dyck is expressible by C-RASP, it only significantly outperforms $k$-Dyck on verbatim retrieval. That being said, $k$-Shuffle Dyck is more efficient than $k$-Dyck, achieving its optimal amount of pre-pretraining 500 steps earlier. Taken together, we find modest evidence supporting the importance of the expressibility of the language in C-RASP.
5 Mechanistic Analysis: Subnetworks
@@
Subnetworks hypothesis: Attention heads established during formal language pre-pretraining are later used to represent the hierarchical structure of natural language.
We test this hypothesis by ablating attention heads of the pre-pretraining subnetwork and comparing the drop in performance against random attention head ablations. Concretely, we pre-pretrain on $\mathcal{D}_{\text{ppt}}$ and prune the model to find the sparse subnetwork $\mathcal{M}(\theta_{t_{0}})$. We use the heuristic core pruning algorithm from Bhaskar et al. (2024), which iteratively removes attention heads from the transformer using structured pruning (Xia et al., 2022) while minimizing the tradeoff between sparsity and language modeling loss on $\mathcal{D}_{\text{ppt}}$. After transfer and training on $\mathcal{D}_{\text{pt}}$, we evaluate the masked model $\mathcal{M}(\theta_{t_{1}})$ against a model $\mathcal{M}_{\text{null}}(\theta_{t_{1}})$ where a subnetwork with the same number of randomly chosen attention heads was ablated.
Positive transfer from $\mathcal{D}_{\text{ppt}}$ to $\mathcal{D}_{\text{pt}}$ could occur for reasons unrelated to subnetworks (e.g., computations are distributed across all heads or in other components of the model). In this case, the masked model $\mathcal{M}$ should perform no better than random masks $\mathcal{M}_{\text{null}}$ when applied to $\theta_{t_{1}}$. However, if pre-pretraining does induce useful inductive biases, we would expect $\mathcal{M}$ to be an important subnetwork even after training on $\mathcal{D}_{\text{pt}}$. So in the alternative hypothesis, $\mathcal{M}$ should significantly outperform $\mathcal{M}_{\text{null}}$.
Results.
After pre-pretraining on $k$-Shuffle Dyck, we ablate $50\%$ of the attention heads. Following previous work Bhaskar et al. (2024); Zhang and Nanda (2024), we replace an ablated head with its mean activation. We compare $\mathcal{M}$ to the random subnetwork $\mathcal{M}_{\text{null}}$ and to $\mathcal{M}$’s complement subnetwork $\mathcal{M}^{c}$.
@@ Results.
We find that $\mathcal{M}$ outperforms $\mathcal{M}_{\text{null}}$ and $\mathcal{M}^{c}$ in both language modeling and grammaticality (Figure 4). We reject the null hypothesis that the subnetwork $\mathcal{M}$ established during pre-pretraining has the same performance as a randomly sampled subnetwork ($p\ll 0.001)$. Further supporting the role of the heads in $\mathcal{M}$, we find that $\mathcal{M}^{c}$, which excludes all of the heads identified by the pruning procedure, performs much more poorly than $\mathcal{M}_{\text{null}}$, that only includes a random subset of them. That being said, while the performance $\mathcal{M}$ is close to the performance of the full network, it does not quite match it, indicating that attention heads outside of $\mathcal{M}$ also play a role in processing natural language.
A breakdown of grammaticality judgment accuracy by grammatical phenomenon shows that only a handful of phenomena are unaffected by masking, some substantially (e.g., the accuracy on subject-verb agreement drops 12 percentage points; see 9 in the Appendix). These phenomena appear to be ones that are syntactically simple but diagnose sensitivity to word structure (morphology), e.g., the distinction between broke and broken; we hypothesize that this aspect of linguistic knowledge is less likely to be mechanistically related to the processing of dependencies in a formal language.
6 Additional Analyses
@@
Figure 5: Blue bars: Testing if the benefit of pre-pretraining on $k$-Shuffle Dyck can be reduced to learning local statistics. We find that pre-pretraining on $n$-gram metamers Kumar et al. (2022) of $k$-Shuffle Dyck performs worse than pre-pretraining on $k$-Shuffle Dyck itself. Purple bar: Vocabulary size manipulation. The best vocabulary size for $k$-Shuffle Dyck is $k=128$.
Figure 6: Pre-pretraining Pythia-1B on 1.6B tokens of $k$-Shuffle Dyck improves over the baselines, especially on language modeling and the retrieval evaluation.
This section reports three additional experiments. Due to the computational cost of pretraining, we focus on $k$-Shuffle Dyck, which performed well in our main experiments. First, we test and rule out the hypothesis that pre-pretraining on $k$-Shuffle Dyck is only effective because of its local statistical properties, and conclude that its effectiveness stems from its structural properties. Next, we study the impact of the vocabulary size hyperparameter $k$ on the effectiveness of transfer from $k$-Shuffle Dyck. Finally, we perform a larger scale training run with Pythia 1B and find that pre-pretraining on $k$-Shuffle Dyck helps in this setting as well. In all of these experiments, we used the optimal number of pre-pretraining gradient updates $t^{*}_{0}=500$ found in our main experiment (equivalent to 30 million tokens).
Transfer is not only due to local statistical properties.
@@
Could the successful transfer from $k$-Shuffle Dyck to natural language be due to the local statistical properties of $k$-Shuffle Dyck, rather than its dependency structure? Learning local statistical regularities is consistent with the finding that neural networks can exhibit distributional simplicity bias (DSB)—they learn simpler statistical patterns, such as the mean and covariance of their representations, before progressing to higher-order relationships Saxe et al. (2014); LeCun et al. (1991); in particular, transformer language models learn $n$-gram statistics in order of increasing complexity Belrose et al. (2024).
@@
To test this hypothesis, we create variants of $k$-Shuffle Dyck that share its local statistics but not its global, rule-based structure. Concretely, we train unigram, bigram, and trigram models on the pre-pretraining corpus we generated from $k$-Shuffle Dyck for our main experiment, and, using these $n$-gram models, we generate “metamer datasets” Kumar et al. (2022) of equivalent size.
We find that pre-pretraining on metamer datasets is strictly less effective than pre-pretraining on $k$-Shuffle Dyck, ruling out the hypothesis that the benefit of pre-pretraining on $k$-Shuffle Dyck is due to local statistics. That being said, pre-pretraining on the unigram metamer performs the worst, followed by bigram and trigram, suggesting that local statistics may explain part of the success of pre-pretraining on structured languages.
Larger vocabulary size may be beneficial.
@@
Pre-pretraining is effective at the 1B scale too.
Finally, we examine whether our results generalize to larger settings by training Pythia-1B on 1.63B tokens from C4 (25,000 steps). In this setting, pre-pretraining on $k$-Shuffle Dyck continues to outperform the baselines on all evaluation metrics (Figure 6) and achieves the final loss of the no-pre-pretraining baseline after training for only 1.10B total tokens. This equates to a token efficiency gain of 33% $\left(1-\frac{1.10\text{B}}{1.63\text{B}}\right)$, or an MRS of 17.3 $\left(\frac{1.63\text{B}-1.10\text{B}}{0.03\text{B}}\right)$. $k$-Shuffle Dyck’s MRS is $\gg$1 for both 160M and 1B training runs, suggesting that pre-pretraining could increase the efficiency of large-scale pretraining as well.
@@ 7 Related Work
The goal of pre-pretraining is similar to that of optimization-based meta-learning, which aims to create a weight initialization that allows the model to rapidly learn new tasks Finn et al. (2017); Nichol et al. (2018) and languages McCoy et al. (2020a); McCoy and Griffiths (2025). The beneficial effect of pre-pretraining on formal language is consistent with the evidence of transfer from source code to natural language, especially for structured tasks Petty et al. (2025); Aryabumi et al. (2025); Kim et al. (2024). In the vision domain, Nakamura et al. (2024) show that a thousand synthetically generated images can replace a million images from ImageNet-1k, in a similar spirit to our work.
Transfer in NLP has also been studied across different languages and domains (Ruder et al., 2019; Pruksachatkun et al., 2020; Deshpande et al., 2022). Most relevant to our work, Mueller and Linzen (2023) show that pretraining on child-directed speech gives a better inductive bias for learning hierarchical syntactic features than standard pretraining corpora. Furthermore, introducing a small amount of synthetic, disambiguating data into pretraining can induce a language model to change its generalization strategy Warstadt et al. (2020b). Related but distinct from our approach are studies that use synthetic data sampled from a formal language to evaluate models’ generalization behavior in a controlled way McCoy et al. (2019); Kim and Linzen (2020); Li et al. (2023).
Pre-pretraining is a form of curriculum learning (Bengio et al., 2009), the approach of actively adjusting properties of the data during training. Recent work has developed algorithms that automate the discovery of language modeling curricula Chen et al. (2025); Jiang et al. (2025), and many language model training recipes introduce different data mixtures at different stages Allal et al. (2025); OLMo et al. (2025); Ouyang et al. (2022). The positive results of our experiments contrast with the largely negative results of the attempts to improve language models’ data efficiency via linguistics-inspired curriculum learning, as part of the BabyLM challenge Warstadt et al. (2023); Hu et al. (2024), pointing to the crucial effect of the particular data presented as part of the curriculum.